how to find the equivalent weight of K4[Fe(CN)6] in the reaction
K4[Fe(CN)6]--->(K^+ ) + (Fe^+3) +CO2+(NO3^-1).If the molar mass of K4[Fe(CN)6] is M

Grade:11

1 Answers

Vikas TU
7 years ago
U need to find the valence factor first for the Fe in K4[Fe(CN)6] .
Oxidation state of Fe is => +2
Valence factor = 3 – 2 = 1
Eq. Weight = M/1 = M